a person has 3 shares in alottery containing 2 prizes and 5 blanks the chance of getting a prize is

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Phani

Chance of getting a prize = chance of getting atleast one prize=1- probability of getting no prize

=1-(5/7)*(4/6)*(3/5)=5/7

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.